Question 4: You roll a fair die once. Define the events A = “the result is one of the numbers 1, 3, and 4”, B = “the result is one of the numbers 3, 4, 5, and 6”. • Are the events A and B independent? Jusitfy your answer using the definition of independence. • Are the events A and B independent? Jusitfy your answer using the definition of conditional probability.

Question 4: You roll a fair die once. Define the events A- B- the result is one of the numbers 1, 3, and 4, the result is one of the numbers 3, 4, 5, and 6 Are the events A and B independent? Jusitfy your answer using the definition of independence. . Are the events A and B independent? Jusitfy your answer using the definition of conditional probability
Answer #1

P(A) = 3/6 = 1/2
P(B) = 4/6 = 2/3

P(A and B) = 2/6 = 1/3

P(A) * P(B) = 1/2 * 2/3 = 1/3

As P(A) * P(B) is equal to P(A and B), events A and B are independent.